You have $50,000 in savings for retirement in an investment earning 12% annually. You aspire to have $1,000,000 in savings when you retire. Assuming you add no more to your savings, how many years will it take to reach your goal? Please round your answer to the nearest hundredth. Note that the HP 12c financial calculator rounds up the periods result to the next integer and will not give the correct answer to the nearest hundredth. Therefore, you should use Excel or a financial calculator that does provide decimal precision to the number of periods.
Answer #1

Answer:
Amount deposited (Present Value) = $50,000
Value of deposit at the time of retirement (Future Value) = $1,000,000
Rate (r ) = 12%
Time (n ) = ??

Future Value = Present Value * (1 + r)^ n
$1,000,000 = $50,000 * (1 + 0.12)^ n
$1,000,000 = $50,000 * (1.12)^ n
$1,000,000 / $50,000 = (1.12)^ n
20 = (1.12)^ n
ln (20) = ln (1.12) * n
n = 26.43 years

It would take 26.43 years to reach goal